Assume the service life is distributed normally, a random sample of several electronic stuff with a service life (years): 6, 7, 7, 7, 6, 8, 8, 7, 6.

Does this indicate that the average lifespan of a shaker from Onix is not the same as 7

year? Do hypothesis testing with a significance level:

a. 1%

b. 5%

c. 10%


1
Expert's answer
2022-05-10T06:43:27-0400
"\bar{x}=\dfrac{6+7+7+7+6+8+8+7+6}{9}"

"=\dfrac{62}{9}\approx6.8889"

"s^2=\dfrac{\sum _i(x_i-\bar{x})^2}{n-1}\approx0.611111"

"s=\sqrt{0.611111}\approx0.7817"

a. The following null and alternative hypotheses need to be tested:

"H_0:\mu=7"

"H_a:\mu\not=7"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=8" degrees of freedom, and the critical value for a two-tailed test is "t_c = 3.355361."

The rejection region for this two-tailed test is "R = \{t: |t| >3.355361\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{6.8889-7}{0.7817/\sqrt{9}}\approx-0.426378"

Since it is observed that "|t|= 0.426378 < 3.355361=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed test, "df=8" degrees of freedom, "t=-0.426378" is "p = 0.681074," and since "p =0.681074 \ge 0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 7, at the "\alpha = 0.01" significance level.


b. The following null and alternative hypotheses need to be tested:

"H_0:\mu=7"

"H_a:\mu\not=7"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=8" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.306002."

The rejection region for this two-tailed test is "R = \{t: |t| >2.306002\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{6.8889-7}{0.7817/\sqrt{9}}\approx-0.426378"

Since it is observed that "|t|= 0.426378 < 2.306002=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed test, "df=8" degrees of freedom, "t=-0.426378" is "p = 0.681074," and since "p =0.681074 \ge 0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 7, at the "\alpha = 0.05" significance level.


c. The following null and alternative hypotheses need to be tested:

"H_0:\mu=7"

"H_a:\mu\not=7"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=8" degrees of freedom, and the critical value for a two-tailed test is "t_c =1.859547."

The rejection region for this two-tailed test is "R = \{t: |t| >1.859547\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{6.8889-7}{0.7817/\sqrt{9}}\approx-0.426378"

Since it is observed that "|t|= 0.426378 < 1.859547=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value for two-tailed test, "df=8" degrees of freedom, "t=-0.426378" is "p = 0.681074," and since "p =0.681074 \ge 0.10=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 7, at the "\alpha = 0.10" significance level.